# Mitigating Long-Tail Bias in HOI Detection via Adaptive Diversity Cache

### v1

@@ Abstract
Human-Object Interaction (HOI) detection is a fundamental task in computer vision, empowering machines to comprehend human-object relationships in diverse real-world scenarios. Recent advances in VLMs have significantly improved HOI detection by leveraging rich cross-modal representations. However, most existing VLM-based approaches rely heavily on additional training or prompt tuning, resulting in substantial computational overhead and limited scalability, particularly in long-tailed scenarios where rare interactions are severely underrepresented. In this paper, we propose the Adaptive Diversity Cache (ADC) module, a novel training-free and plug-and-play mechanism designed to mitigate long-tail bias in HOI detection. ADC constructs class-specific caches that accumulate high-confidence and diverse feature representations during inference. The method incorporates frequency-aware cache adaptation that favors rare categories and is designed to enable robust prediction calibration without requiring additional training or fine-tuning. Extensive experiments on HICO-DET and V-COCO datasets show that ADC consistently improves existing HOI detectors, achieving up to +8.57% mAP gain on rare categories and +4.39% on the full dataset, demonstrating its effectiveness in mitigating long-tail bias while preserving overall performance.
1 Introduction
@@
Human-Object Interaction (HOI) detection [gao2018ican] stands as a cornerstone in computer vision, with broad applications in autonomous driving and robotic manipulation [jia2025contexthoi, liu2022interactiveness, wu2024toward, gao2024dual, guo2024unseen, lei2024ez, li2025cogvla]. By identifying interactions between humans and objects, such as “holding a bottle” or “riding a bicycle”, HOI detection delivers enriched semantic understanding to facilitate advanced decision-making. Moreover, HOI detection serves as a foundational component for tasks like action recognition [jhuang2013towards, sun2022human, li2025prototypical, li2025frame] and visual question answering [antol2015vqa, yang2025multimodal, cocchi2025augmenting, li2025lion].
@@
Vision-Language Models (VLMs), such as CLIP [radford2021learning] and Flamingo [alayrac2022flamingo], achieve robust alignment between visual and language modalities through pre-training on large-scale image-text pairs. These models excel in zero-shot and few-shot learning tasks, effectively bridging the gap between visual features and semantic descriptions to provide rich cross-modal representations for downstream applications. As illustrated in Fig. 1(a) and (b), existing VLM-based HOI detection approaches can be broadly categorized into two paradigms: alignment-based methods and prompt-based methods. Alignment-based methods integrate HOI-related visual features with pre-trained VLMs through feature alignment to exploit implicit knowledge for interaction recognition. These approaches typically employ knowledge distillation or feature optimization techniques to improve the alignment between visual and textual representations. In contrast, prompt-based methods adapt VLMs to HOI tasks via prompt learning without requiring major architectural modifications. Such methods introduce trainable prompts for multi-modal optimization and leverage guided prompt learning strategies to enhance model performance.
@@
Despite the remarkable progress achieved in HOI recognition, existing methods remain reliant on additional training procedures. This reliance not only incurs substantial computational costs and the demand for large-scale annotated data, but also limits the scalability and deployment of HOI systems in real-world scenarios where annotation resources are scarce or unavailable. Additionally, these training-based approaches often struggle to generalize to rare interactions, particularly under the long-tailed distribution prevalent in HOI datasets (see Fig. 2). A handful of frequent interactions dominate in such distributions, while rare HOIs are severely underrepresented. This imbalance induces significant prediction bias: models overfit to common categories during training, resulting in insufficient representations that inadequately capture the distinctive features of rare categories. Consequently, recognition performance on rare interactions deteriorates, severely limiting generalization to open-world scenarios.
@@
2 Related Works
2.1 Human-Object Interaction Detection
@@
HOI detection aims to recognize interactions between humans and objects in visual scenes, addressing challenges in tasks like scene understanding and action recognition. HOI methods based on VLMs fall into two paradigms: alignment-based and prompt-based approaches. Alignment-based methods [liao2022gen, ning2023hoiclip, mao2023clip4hoi, cao2023detecting] align HOI visual features with pre-trained VLMs to leverage their knowledge for interaction detection. Representative works include GEN-VLKT [liao2022gen], which transfers CLIP [radford2021learning] semantic representations through knowledge distillation, and HOICLIP [ning2023hoiclip], which directly utilizes VLM visual encoders for HOI feature extraction. Prompt-based approaches modify VLMs through prompt learning for HOI tasks [khattak2023maple, guo2024unseen, lei2024ez]. Notable examples are MaPLe [khattak2023maple], which employs multi-modal prompt tuning, and EZ-HOI [lei2024ez], which introduces guided prompt learning with foundation model guidance to enhance performance. However, these methods often require additional training or fine-tuning phases, rely on domain-specific designs, and are susceptible to noise in continuous data streams, leading to slow adaptation convergence.
2.2 Test Time Adaptation
@@
Test-time adaptation (TTA) [wang2020tent, qiao2025bidirectional] aims at updating a trained model on a single unlabeled test sample before making predictions to increase its robustness to distribution shift. Early TTA work primarily focused on image classification, with methods like TENT [wang2020tent] minimizing entropy of model predictions, and others exploring self-supervised auxiliary tasks [sun2020test] and contrastive learning approaches [chen2022contrastive]. Recent advances include TDA [karmanov2024efficient], which introduces a training-free dynamic adapter with key-value cache for efficient adaptation, and DPE [zhang2024dual], which evolves dual prototypes to accumulate task-specific knowledge from multi-modalities in VLMs. The success in classification has inspired applications to semantic segmentation [hu2021fully, shin2022mm], image super-resolution [deng2023efficient, shocher2018zero], and object detection [ruan2024fully, veksler2023test]. However, HOI detection [ning2023hoiclip, lei2024ez] presents unique challenges due to its compositional nature, where the same action can involve different objects and vice versa, making it particularly susceptible to domain shifts. Introducing TTA to HOI detection faces challenges, including complex pseudo-label generation for multi-faceted predictions, difficult confidence estimation for spatial-semantic relationships, and adaptation to long-tail HOI distributions. Despite these challenges, we present a novel plug-and-play module for test-time adaptation in HOI detection, aiming to address the long-tail problem and enable robust performance across diverse interaction patterns.
3 Method
@@
This section introduces the Adaptive Diversity Cache (ADC) mechanism, a novel training-free and plug-and-play module designed to mitigate the long-tail bias in HOI detection. Unlike existing TTA methods (e.g., BoostAdapter [zhang2024boostadapter]) that focus on parameter adaptation or feature distribution alignment for single-label classification, our ADC module addresses the compositional nature of HOI detection. While TDA employs key-value caches for feature adaptation, it lacks mechanisms to handle the severe class imbalance(long-tailed distribution) and compositional complexity inherent in HOI tasks.
@@
3.1 Confidence-Diversity Joint Cache Selection
Inspired by Boostadapter [zhang2024boostadapter], we posit that maintaining a historical prediction queue of HOI pair visual features alleviates inference bias from imbalanced data. This dynamic repository enables reference to diverse historical information, counteracting skewed learning from uneven class distribution.
@@ 3.1 Confidence-Diversity Joint Cache Selection
The feature diversity score is defined via multi-scale geometric analysis, capturing both local and global distributional properties:
@@
Fig. 4 presents t-SNE [hinton2008visualizing] visualizations of features stored in () on HICO-DET after 1000 (left) and 8000 (right) updates. Highlighting 25 random classes, the plots demonstrate progressive refinement of CJCS representativeness: features form tighter clusters over time, with semantically similar interactions clustering coherently, validating the cache’s ability to capture discriminative HOI patterns. Note that verb-sharing interactions (with distinct objects) naturally co-cluster due to shared verb semantics, which does not impair performance as object priors resolve ambiguities during inference.
3.2 Frequency-Aware Cache Adaptation
@@
4 Experiments
4.1 Datasets and Metrics
4.1.1 Datasets
We conduct experiments on two widely-used HOI detection datasets, HICO-DET [chao2018learning] and V-COCO [gupta2015visual]. The HICO-DET dataset is a large-scale benchmark for HOI detection. It defines 600 interaction types as combinations of 80 object categories and 117 action classes. The dataset comprises 38,118 training images and 9,658 test images, supporting comprehensive model training and evaluation. In contrast to HICO-DET, the V-COCO dataset is much smaller in scale, containing 5,400 train-val images and 4,946 test images. It includes 80 object categories, 29 action classes, and 263 interaction types. V-COCO offers complementary evaluation scenarios with distinct interaction complexities and distribution characteristics.
4.1.2 Metrics
We adopt the mean Average Precision (mAP) as the primary evaluation metric, following [chao2018learning]. For HICO-DET, we report mAP scores with three different category configurations: (1) Full: evaluation on all 600 interaction types; (2) Rare: evaluation on long-tail categories with fewer than 10 training samples; (3) Non-rare: evaluation on the remaining categories with sufficient training data. For V-COCO, we report the role mAP under the Scene 1 setting, which evaluates performance across all interaction types following the standard evaluation protocol.
4.2 Implementation Details
We employ DETR [carion2020end] as the object detector, which has been pre-trained on HICO-DET and V-COCO datasets with a ResNet-50 backbone. We adopt ADA-CM [lei2023efficient] as the baseline model and freeze all network parameters. In test-time adaptation, we first feed the entire test set into the model to obtain a complete cache initialization, and subsequently utilize the established cache to generate final prediction results while continuing to update the cache using the enhanced predictions. In the FACA module, we apply random feature augmentation operations, including crop, rotate, shear, and translate, to both human and object features, generating 64 augmented features for each sample and selecting the top 10% with the lowest self-entropy values. The augmentation intensity is controlled by a severity parameter, which is set to 1 by default, and the shot capacity of the cache is configured to 6. All experiments are conducted with a single Tesla A100 40GB GPU. It’s worth noting that ADC can be seamlessly integrated into different architectural paradigms without modification, as demonstrated by our experiments across multiple baseline models. The only requirement is that the base detector outputs interaction logits during inference, making it applicable to virtually all existing HOI methods.
@@ 4.2 Implementation Details
| iCAN [gao2018ican] | R50 | 14.84 | 10.45 | 16.15 | 45.3 |
| DRG [gao2020drg] | R50-F | 24.53 | 19.47 | 26.04 | 51.0 |
| HOTR [kim2021hotr] | R50 | 25.10 | 17.34 | 27.42 | 55.2 |
| QPIC [tamura2021qpic] | R50 | 29.07 | 21.85 | 31.23 | 58.8 |
| CPC [park2022consistency] | R50 | 29.63 | 23.14 | 31.57 | 63.1 |
| FGAHOI [ma2023fgahoi] | Swin-T | 29.94 | 22.24 | 32.24 | 60.5 |
| RLIP [yuan2022rlip] | R50 | 32.84 | 26.85 | 34.63 | 61.9 |
| UPT [zhang2022efficient] | R50 | 32.62 | 28.62 | 33.81 | 59.0 |
| GENVLKT [liao2022gen] | R50 | 33.75 | 29.25 | 35.10 | 62.4 |
| PViC [zhang2023exploring] | R50 | 34.69 | 32.14 | 35.45 | 62.8 |
| HOICLIP [ning2023hoiclip] | R50 | 34.69 | 31.12 | 35.74 | 63.5 |
| RLIPv2 [yuan2023rlipv2] | R50 | 35.38 | 29.61 | 37.10 | 65.9 |
| LOGICHOI [li2023neural] | R50 | 35.47 | 32.03 | 36.22 | 64.4 |
| ViPLO [park2023viplo] | ViT-B | 37.22 | 35.45 | 37.75 | 62.2 |
| SCTC [jiang2024exploring] | R50 | 37.92 | 34.78 | 38.86 | 67.1 |
| HOIGEN [guo2024unseen] | R50 | 33.56 | 32.10 | 34.00 | - |
| + BoostAdapter [zhang2024boostadapter] | R50 | 34.83 | 35.37 | 34.47 | - |
@@
4.3 Comparison with State-of-the-Art
Tab. 1 presents performance comparisons of our ADC module integrated with different baseline methods on HICO-DET and V-COCO datasets. When equipped with the ADC module, ADA-CM with a ResNet-50 backbone achieves 39.81 full mAP and an impressive 41.48 mAP on the rare split, establishing new state-of-the-art performance. Compared to BoostAdapter, a representative TTA method, ADC demonstrates superior performance gains: while BoostAdapter achieves modest improvements (+0.36 mAP full, +1.55 mAP rare for ADA-CM), our method delivers substantially larger gains (+1.41 mAP full, +3.96 mAP rare). This indicates that our approach successfully addresses the fundamental challenge of long-tail distribution without sacrificing performance on non-rare categories. Furthermore, the plug-and-play nature of our ADC module is validated by its consistent effectiveness across different baseline methods. As shown in Tab. 1, ADC achieves improvements when integrated with HOIGEN (+8.57 mAP on rare categories) and EZ-HOI (+5.15 mAP on rare categories), demonstrating its broad applicability and generalization capability.

### v2

@@ Abstract
Human-Object Interaction (HOI) detection is a fundamental task in computer vision, empowering machines to comprehend human-object relationships in diverse real-world scenarios. Recent advances in VLMs have significantly improved HOI detection by leveraging rich cross-modal representations. However, most existing VLM-based approaches rely heavily on additional training or prompt tuning, resulting in substantial computational overhead and limited scalability, particularly in long-tailed scenarios where rare interactions are severely underrepresented. In this paper, we propose the Adaptive Diversity Cache (ADC) module, a novel training-free and plug-and-play mechanism designed to mitigate long-tail bias in HOI detection. ADC constructs class-specific caches that accumulate high-confidence and diverse feature representations during inference. The method incorporates adaptive capacity allocation favoring rare categories and dynamic feature augmentation to enable robust prediction calibration without requiring additional training or fine-tuning. Extensive experiments on HICO-DET and V-COCO datasets show that ADC consistently improves existing HOI detectors, particularly enhancing rare category detection while preserving overall performance. These findings confirm the effectiveness of ADC as a training-free, plug-and-play solution for long-tail bias mitigation.
I Introduction
@@
Human-Object Interaction (HOI) detection [13] stands as a cornerstone in computer vision, with broad applications in autonomous driving and robotic manipulation [11, 7, 21, 35, 54, 14, 15, 26, 32]. By identifying interactions between humans and objects, such as “holding a bottle” or “riding a bicycle”, HOI detection delivers enriched semantic understanding to facilitate advanced decision-making. Moreover, HOI detection serves as a foundational component for tasks like action recognition [33, 20, 48, 29, 28] and visual question answering [51, 2, 56, 9, 31].
@@
Vision-Language Models (VLMs), such as CLIP [43] and Flamingo [1], achieve robust alignment between visual and language modalities through pre-training on large-scale image-text pairs. These models excel in zero-shot and few-shot learning tasks, effectively bridging the gap between visual features and semantic descriptions to provide rich cross-modal representations for downstream applications. As illustrated in Fig. 1(a) and (b), existing VLM-based HOI detection approaches can be broadly categorized into two paradigms: alignment-based methods and prompt-based methods. Alignment-based methods integrate HOI-related visual features with pre-trained VLMs through feature alignment to exploit implicit knowledge for interaction recognition. These approaches typically employ knowledge distillation or feature optimization techniques to improve the alignment between visual and textual representations. In contrast, prompt-based methods adapt VLMs to HOI tasks via prompt learning without requiring major architectural modifications. Such methods introduce trainable prompts for multi-modal optimization and leverage guided prompt learning strategies to enhance model performance.
@@
Despite the remarkable progress achieved in HOI recognition, existing methods remain reliant on additional training procedures. This reliance not only incurs substantial computational costs and the demand for large-scale annotated data, but also limits the scalability and deployment of HOI systems in real-world scenarios where annotation resources are scarce or unavailable. Additionally, these training-based approaches often struggle to generalize to rare interactions, particularly under the long-tailed distribution prevalent in HOI datasets (see Fig. 2). A handful of frequent interactions dominate in such distributions, while rare HOIs are severely underrepresented. This imbalance induces significant prediction bias: models overfit to common categories during training, resulting in insufficient representations that inadequately capture the distinctive features of rare categories. It is important to note that the long-tail issue in HOI detection is multi-factorial. It stems not only from the frequency imbalance of object or verb categories but fundamentally from the compositional sparsity of HOI triplets [18]. Many valid human-object combinations appear rarely or never in the training set, making it difficult for models to learn robust visual prototypes for these ”tail” compositions. Consequently, recognition performance on rare interactions deteriorates, severely limiting generalization to open-world scenarios.
@@
II Related Works
II-A Human-Object Interaction Detection
@@
HOI detection aims to recognize interactions between humans and objects in visual scenes, addressing challenges in tasks like scene understanding and action recognition. HOI methods based on VLMs fall into two paradigms: alignment-based and prompt-based approaches. Alignment-based methods [34, 38, 37, 3] align HOI visual features with pre-trained VLMs [55, 8, 41] to leverage their knowledge for interaction detection. Representative works include GEN-VLKT [34], which transfers CLIP [43] semantic representations through knowledge distillation, and HOICLIP [38], which directly utilizes VLM visual encoders for HOI feature extraction. Prompt-based approaches modify VLMs through prompt learning for HOI tasks [24, 15, 26]. Notable examples are MaPLe [24], which employs multi-modal prompt tuning, and EZ-HOI [26], which introduces guided prompt learning with foundation model guidance to enhance performance. However, these methods often require additional training or fine-tuning phases, rely on domain-specific designs, and are susceptible to noise in continuous data streams, leading to slow adaptation convergence.
II-B Test Time Adaptation
@@
Test-time adaptation (TTA) [53, 52, 42] aims to update a trained model on a single unlabeled test sample before making predictions to increase its robustness to distribution shift. Early TTA work primarily focused on image classification, with methods like TENT [52] minimizing entropy of model predictions, and others exploring self-supervised auxiliary tasks [47] and contrastive learning approaches [6]. Recent advances include TDA [23], which introduces a training-free dynamic adapter with key-value cache for efficient adaptation, and DPE [59], which evolves dual prototypes to accumulate task-specific knowledge from multi-modalities in VLMs. The success in classification has inspired applications to semantic segmentation [19, 45], image super-resolution [10, 46], and object detection [44, 50]. However, HOI detection [38, 26] presents unique challenges due to its compositional nature, where the same action can involve different objects and vice versa, making it particularly susceptible to domain shifts. Introducing TTA to HOI detection faces challenges, including complex pseudo-label generation for multi-faceted predictions, difficult confidence estimation for spatial-semantic relationships, and adaptation to long-tail HOI distributions. Despite these challenges, we present a novel plug-and-play module for test-time adaptation in HOI detection, aiming to address the long-tail problem and enable robust performance across diverse interaction patterns.
III Method
@@
This section introduces the Adaptive Diversity Cache (ADC) mechanism, a novel training-free and plug-and-play module designed to mitigate the long-tail bias in HOI detection. Unlike existing TTA methods (e.g., BoostAdapter [62]) that focus on parameter adaptation or feature distribution alignment for single-label classification, our ADC module addresses the compositional nature of HOI detection. While TDA employs key-value caches for feature adaptation, it lacks mechanisms to handle the severe class imbalance(long-tailed distribution) and compositional complexity inherent in HOI tasks.
@@
III-A Confidence-Diversity Joint Cache Selection
Inspired by Boostadapter [62], we posit that maintaining a historical prediction queue of HOI pair visual features alleviates inference bias from imbalanced data. This dynamic repository enables reference to diverse historical information, counteracting skewed learning from uneven class distribution.
@@ III-A Confidence-Diversity Joint Cache Selection
The feature diversity score is defined via multi-scale geometric analysis, capturing both local and global distributional properties, where ”local” refers to neighborhood sensitivity (), and ”global” denotes angular diversity relative to the cache:
@@
Fig. 4 presents t-SNE [17] visualizations of features stored in () on HICO-DET after 1000 (left) and 8000 (right) updates. Highlighting 25 random classes, the plots demonstrate progressive refinement of CJCS representativeness: features form tighter clusters over time, with semantically similar interactions clustering coherently, validating the cache’s ability to capture discriminative HOI patterns. Note that verb-sharing interactions (with distinct objects) naturally co-cluster due to shared verb semantics, which does not impair performance as object priors resolve ambiguities during inference.
III-B Frequency-Aware Cache Adaptation
@@
IV Experiments
IV-A Datasets and Metrics
IV-A1 Datasets
We conduct experiments on two widely-used HOI detection datasets, HICO-DET [5] and V-COCO [16]. The HICO-DET dataset is a large-scale benchmark for HOI detection. It defines 600 interaction types as combinations of 80 object categories and 117 action classes. The dataset comprises 38,118 training images and 9,658 test images, supporting comprehensive model training and evaluation. In contrast to HICO-DET, the V-COCO dataset is much smaller in scale, containing 5,400 train-val images and 4,946 test images. It includes 80 object categories, 29 action classes, and 263 interaction types. V-COCO offers complementary evaluation scenarios with distinct interaction complexities and distribution characteristics.
IV-A2 Metrics
We adopt the mean Average Precision (mAP) as the primary evaluation metric, following [5]. For HICO-DET, we report mAP scores with three different category configurations: (1) Full: evaluation on all 600 interaction types; (2) Rare: evaluation on long-tail categories with fewer than 10 training samples; (3) Non-rare: evaluation on the remaining categories with sufficient training data. For V-COCO, we report the role mAP under the Scene 1 setting, which evaluates performance across all interaction types following the standard evaluation protocol.
IV-B Implementation Details
We employ DETR [4] as the object detector, which has been pre-trained on HICO-DET and V-COCO datasets with a ResNet-50 backbone. We adopt ADA-CM [27] as the baseline model and freeze all network parameters. In test-time adaptation, we first feed the entire test set into the model to obtain a complete cache initialization, and subsequently utilize the established cache to generate final prediction results while continuing to update the cache using the enhanced predictions. In the FACA module, we apply random feature augmentation operations, including crop, rotate, shear, and translate, to both human and object features, generating 64 augmented features for each sample and selecting the top 10% with the lowest self-entropy values. The augmentation intensity is controlled by a severity parameter, which is set to 1 by default, and the shot capacity of the cache is configured to 6. All experiments are conducted with a single Tesla A100 40GB GPU. It is worth noting that ADC can be seamlessly integrated into different architectural paradigms without modification, as demonstrated by our experiments across multiple baseline models. The only requirement is that the base detector outputs interaction logits during inference, making it applicable to virtually all existing HOI methods.
@@ IV-B Implementation Details
| iCAN [13] | R50 | 14.84 | 10.45 | 16.15 | 45.3 |
| DRG [12] | R50-F | 24.53 | 19.47 | 26.04 | 51.0 |
| HOTR [25] | R50 | 25.10 | 17.34 | 27.42 | 55.2 |
| QPIC [49] | R50 | 29.07 | 21.85 | 31.23 | 58.8 |
| CPC [40] | R50 | 29.63 | 23.14 | 31.57 | 63.1 |
| FGAHOI [36] | Swin-T | 29.94 | 22.24 | 32.24 | 60.5 |
| RLIP [57] | R50 | 32.84 | 26.85 | 34.63 | 61.9 |
| UPT [60] | R50 | 32.62 | 28.62 | 33.81 | 59.0 |
| GENVLKT [34] | R50 | 33.75 | 29.25 | 35.10 | 62.4 |
| PViC [61] | R50 | 34.69 | 32.14 | 35.45 | 62.8 |
| HOICLIP [38] | R50 | 34.69 | 31.12 | 35.74 | 63.5 |
| RLIPv2 [58] | R50 | 35.38 | 29.61 | 37.10 | 65.9 |
| LOGICHOI [30] | R50 | 35.47 | 32.03 | 36.22 | 64.4 |
| ViPLO [39] | ViT-B | 37.22 | 35.45 | 37.75 | 62.2 |
| SCTC [22] | R50 | 37.92 | 34.78 | 38.86 | 67.1 |
| HOIGEN [15] | R50 | 33.56 | 32.10 | 34.00 | - |
| + BoostAdapter [62] | R50 | 34.83 | 35.37 | 34.47 | - |
@@
IV-C Comparison with State-of-the-Art
Tab. I presents performance comparisons of our ADC module integrated with different baseline methods on HICO-DET and V-COCO datasets. When equipped with the ADC module, ADA-CM with a ResNet-50 backbone achieves 39.81 full mAP and an impressive 41.48 mAP on the rare split, establishing new state-of-the-art performance. Compared to BoostAdapter, a representative TTA method, ADC demonstrates superior performance gains: while BoostAdapter achieves modest improvements (+0.36 mAP full, +1.55 mAP rare for ADA-CM), our method delivers substantially larger gains (+1.41 mAP full, +3.96 mAP rare). This indicates that our approach successfully addresses the fundamental challenge of long-tail distribution without sacrificing performance on non-rare categories. Furthermore, the plug-and-play nature of our ADC module is validated by its consistent effectiveness across different baseline methods. As shown in Tab. I, ADC achieves improvements when integrated with HOIGEN (+8.57 mAP on rare categories) and EZ-HOI (+5.15 mAP on rare categories), demonstrating its broad applicability and generalization capability.